OpenStudy (anonymous):

Perform the modular arithmetic operation. [(11 ∙ 4. + 6] (mod 8) A. 1 B. 5 C. 3 D. 2

OpenStudy (anonymous):

You can either do the arithmetic like normal, then reduce your answer mod 8, or reduce everything mod 8, then do the arithmetic. It doesnt matter.

OpenStudy (campbell_st):

11 x 4 + 6 = 50 50/ 8 = 6 remainder 2 so its mod 2

OpenStudy (anonymous):

So you could do 11*4+6=50, and 50 mod 8 is 2, or reduce 11 mod 8, which is 3, times 3 is 12, reduced mod 8 is 4, plus 6 is 10, reduced mod 8 is 2.

OpenStudy (anonymous):

times 4, not 3, my bad.

OpenStudy (anonymous):

maybe it is \(\equiv 2(\text{ mod 8})\)